# BLAST: A Stealthy Backdoor Leverage Attack against Cooperative Multi-Agent Deep Reinforcement Learning based Systems

### v1

@@ Abstract
Recent studies have shown that cooperative multi-agent deep reinforcement learning (c-MADRL) is under the threat of backdoor attacks. Once a backdoor trigger is observed, it will perform malicious actions leading to failures or malicious goals. However, existing backdoor attacks suffer from several issues, e.g., instant trigger patterns lack stealthiness, the backdoor is trained or activated by an additional network, or all agents are backdoored. To this end, in this paper, we propose a novel Backdoor Leverage Attack againST c-MADRL, BLAST, which attacks the entire multi-agent team by embedding the backdoor only in a single agent. Firstly, we introduce adversary spatiotemporal behavior patterns as the backdoor trigger rather than manual-injected fixed visual patterns or instant status and control the period to perform malicious actions. This method can guarantee the stealthiness and practicality of BLAST. Secondly, we hack the original reward function of the backdoor agent via unilateral guidance to inject BLAST, so as to achieve the leverage attack effect that can pry open the entire multi-agent system via a single backdoor agent. We evaluate our BLAST against 3 classic c-MADRL algorithms (VDN, QMIX, and MAPPO) in 2 popular c-MADRL environments (SMAC and Pursuit), and 2 existing defense mechanisms. The experimental results demonstrate that BLAST can achieve a high attack success rate while maintaining a low clean performance variance rate.
@@ I Introduction
Cooperative multi-agent deep reinforcement learning (c-MADRL) is a significant branch of deep reinforcement learning (DRL). c-MADRL enables multiple agents to cooperate to achieve the same goal to solve complex tasks [1] and has found applications in many areas such as cooperative game [2], autonomous driving [3], computation offloading [4], and other fields [1]. However, current research [5, 6, 7, 8, 9, 10, 11] has revealed that DRL is vulnerable to a serious threat known as backdoor attacks. This attack involves injecting malicious triggers into the training data or the DRL model. Once a DRL model is injected with a backdoor, it will behave in an unexpected or even malicious way when the trigger occurs. Since c-MADRL inherits the characteristics of DRL, it also suffers from this threat.
@@
Figure 1: A comparison of backdoor attacks against c-MADRL. (a) Most existing attacks use instant visual pattern triggers and require the implantation of backdoors into all agents to achieve the attack effect. (b) Our BLAST uses spatiotemporal behavior pattern triggers and has leverage effects that only require implanting the backdoor into a single agent but can achieve the system-wide attack effect.
@@
As of now, only a handful of studies [12, 13, 14, 15] have delved into backdoor attacks against c-MADRL. Attacks [13, 12] directly implant backdoors in all agents and trigger them multiple times to carry out the attack, as shown in Fig. 1(a). While it is rather effortless to implant backdoor attacks in all agents, this approach comes with a high injection cost and lacks sufficient concealment. Attacks [14, 15] poison a single agent to attack the entire system, but ignore the mutual influence between agents [16]. This influence is the key to achieving fast cooperation in multi-agent systems [16] and can therefore weaken or even block the effectiveness of attacks [17]. In addition, existing attack methods all use specific instant visual patterns (as shown in Fig. 1(a)) as triggers, e.g., visual modified map patterns [13], spectrum signals [12], and agent spatial distance [15]. Once the trigger is observed, they immediately execute instantaneous or continuous attack actions. However, this paired backdoor triggering and attack behavior can be easily detected by data anomaly detectors, e.g., [18, 19]. Hence, although these backdoor attacks are effective against c-MADRL, they are still far from perfect. We believe that a perfect one should possess two key characteristics: 1) good concealment and operability, and 2) minimal injection cost with a high impact.
@@
To achieve these two characteristics, we propose a novel stealthy backdoor leverage attack against c-MADRL, BLAST, which only requires implanting a backdoor in a single agent to pry an entire multi-agent team into failure, as shown in Fig. 1(b). We call this effect the leverage attack effect. First, most existing c-MADRL algorithms (e.g., VDN [20], QMIX [21], COMA [22], and MAPPO [23]) use recurrent neural networks (RNN) to memorize past information and make effective decisions in combination with current observations, thereby overcoming partial observability [24]. This, however, also allows attackers to hide backdoor triggers in unobservable states [25]. Hence, to achieve the first characteristic, we distribute our BLAST’s trigger across a sequence of observations over a short period rather than within a single epoch. Once BLAST is deployed, the attacker only needs to act as a moving object (e.g., a user-controlled enemy unit in StarCraft) and perform a specific sequence of actions around the BLAST agent [^1] In the follow-up, we use this term to denote a c-MADRL agent that is implanted with our BLAST backdoor attack. in a short period following the trigger to activate the attack. Besides, we perform BLAST’s attack behaviors within a controllable period after the trigger occurs, rather than immediately. This method of decoupling attack triggers and actions and distributing them in different time series can ensure that BLAST can achieve high concealment and a high attack success rate at a low poisoning rate.
@@
For the second characteristic, BLAST does not require additional networks to achieve the leverage attack effect. Specifically, c-MADRL agents are commonly trained by enlarging mutual influence between agents to achieve faster cooperation [16]. Under this influence, actions taken by agents will affect each other. Based on this, we design a reward hacking method for backdoor injection, in which we introduce the unilateral influence filter [17] to only enlarge the influence of the BLAST agent on other agents, but eliminate the opposite influence. It consists of two reward items: the former encourages the BLAST agent to perform actions that have long-term malicious effects on clean agents through target failure state guidance; the latter incentivizes the BLAST agent to quickly find actions that can mislead its teammates into performing non-optimal actions. With this hacking method, BLAST can quickly attack the entire system by poisoning only one agent.
@@
  • We use spatiotemporal behavior patterns as triggers of our BLAST and its attack period is controllable, which enhances its concealment and operability;
@@ II Background
Most of the existing popular c-MADRL algorithms adopt the centralized training and decentralized execution (CTDE) framework [27, 28]. The CTDE framework centralizes the evaluation of the joint policies of all agents during training to address the issue of environmental non-stationarity. During execution, each agent is dispersed and executed separately, relying solely on its own local observation and observation history for action selection, greatly alleviating the problem of low efficiency in policy execution. c-MADRL can be roughly divided into two categories: value function decomposition (VFD) and centralized value function (CVF).
@@
CVF-based c-MADRL typically addresses the scalability and environmental non-stationarity issues by learning a centralized Critic network and each independent Actor network for each agent. MAPPO [23] is a CVF-based algorithm that applies the PPO algorithm [29] to multi-agent environments. Through methods such as value function normalization, action masking, and importance sampling, it effectively solves the problem of low data sample sampling efficiency in on-policy algorithms under limited computing resources. The objective function of MAPPO is as follows:
@@ III-B Attacker’s Capacities and Goals
Attacker’s Goals. With the above capabilities and limitations, the attacker’s main goal is to attack the entire multi-agent team by triggering the backdoor in a single agent during execution. Moreover, the implanted backdoor needs to be effective and stealthy. Specifically, if the backdoor trigger is present, the BLAST agent is capable of behaving maliciously or abnormally to disrupt the entire team, otherwise it is capable of behaving normally like a clean agent. Besides, the backdoor triggers should be concealed, occur as infrequently as possible, and have a low poisoning rate.
@@ IV-A Spatiotemporal Backdoor Trigger
IV-B Reward Hacking based Unilateral Influence
In c-MADRL systems, all agents typically interact with each other, thus there is a mutual influence relationship between agents. Due to this mutual influence, the sphere of influence of reward hacking methods in the existing backdoor attacks [5, 6, 7, 8, 10, 11] may be limited only to the backdoor agent itself rather than the entire system during the attack period. Therefore, we design a new reward hacking method based on the unilateral influence filter [17] which can eliminate the detrimental influence from other agents to the backdoor agent and only enable the influence from the backdoor agent to other agents.
@@ IV-B Reward Hacking based Unilateral Influence
To enhance the effectiveness of BLAST, we further consider from the perspective of action deviation, where the BLAST agent performs a malicious action causing the clean agents’ actions at the next time step to deviate from the original optimal actions. We set the hacked reward term based on action deviation as follows:
@@ IV-C BLAST Model Training
Firstly, the attacker specifies a spatiotemporal behavior trigger $\mathcal{T}:=(\Psi,\zeta)$ with a trigger period $\mathbb{N}^{t}$. Besides, an attack period $L$ is introduced to attack only $L$ time steps after the trigger appears completely, which can enhance the stealthiness of the backdoor attacks. In each training episode, the attacker determines whether to poison with the poisoning rate $p$. If poison, the attacker will insert the backdoor trigger (Line 3). At each time step $t$ in an episode, clean agents select their actions according to their policies, and the BLAST agent chooses a random action with the probability $\sigma$, otherwise, it chooses an action according to its policy $\pi^{b}$ (Line 5-7). During the attack period $L$, the attacker will hack the reward $r_{t}$ as Equ (5) (Line 11-14). To ensure the effective training of the BLAST model, we set up two replay buffers, $\mathcal{B}_{p}$ and $\mathcal{B}_{c}$, to store poisoned and clean episodes, respectively (Line 17-21). To update the BLAST model, we randomly sample a batch of episode data from replay buffer $\mathcal{B}_{p}$ or $\mathcal{B}_{c}$ with probability $p$ or $1-p$ (Line 22-23).
@@ V-A1 Environments
We use StarCraft Multi-Agent Challenge (SMAC) [2] and Pursuit [30] in pettingzoo [31] as experimental environments, as shown in Fig. 3: (a) SMAC is an environment of two teams against each other where all allied units (i.e., the red side) are controlled by c-MADRL agents, while enemy units (i.e., the blue side) are controlled by computer. If the enemies’ health decreases, the agent team will receive corresponding positive rewards; If an enemy unit dies, the agent team will receive a reward of 10; If all enemy units die (i.e. win), the agent team will receive a reward of 200; If all allied units die (i.e. fail), the agent team will receive a reward of 0. We choose 8m, 3m, and 2s3z as our test maps, with 8, 3, and 5 agents in the agent team, respectively. (b) In the Pursuit environment, the pursuers are controlled by agents. Every time the agents fully surround an evader, each of the surrounding agents receives a reward of 5, and the evader is removed from the environment. Agents also receive a reward of 0.01 every time they touch an evader. We set the number of agents to 8 and the number of evaders to 10. Each agent receives a fixed penalty of -0.01 per time step, which means that the longer it takes to catch the same number of evaders, the lower the team reward.
@@ V-A3 Backdoor Trigger and Attack Settings
We use the trigger1 $\mathcal{T}_{1}:=(\Psi_{1},\zeta_{1})$ defined in Example 1 to evaluate our backdoor attacks in 8m and 2s3z map. In the 3m map and Pursuit environment, we use trigger 2 and trigger 3 defined as follows, respectively:
  • Trigger 2 $\mathcal{T}_{2}:=(\Psi_{2},\zeta_{2})$, where $\Psi_{2}(\mathbb{N}^{t}=4):=\wedge\psi_{(t-3,x)}\wedge\psi_{(t-3,y)}\wedge\psi_{(t-2,x)}\wedge\psi_{(t-2,y)}\wedge\psi_{(t-1,x)}\wedge\psi_{(t-1,y)}\wedge\psi_{(t,x)}\wedge\psi_{(t,y)}$, and $\zeta_{2}:=(a^{e}_{t-3},a^{e}_{t-2},a^{e}_{t-1},a^{e}_{t})$, where $\psi_{(t-3,x)}:=0.98<x^{e}-x^{b}<1.00$, $\psi_{(t-3,y)}:=-0.10<y^{e}-y^{b}<0.10$, $\psi_{(t-2,x)}:=0.60<x^{e}-x^{b}<0.62$, $\psi_{(t-2,y)}:=0.60<y^{e}-y^{b}<0.62$, $\psi_{(t-1,x)}:=0.60<x^{e}-x^{b}<0.62$, $\psi_{(t-1,y)}:=-0.62<y^{e}-y^{b}<-0.60$, $\psi_{(t,x)}:=0.98<x^{e}-x^{b}<1.00$, $\psi_{(t,y)}:=-0.01<y^{e}-y^{b}<0.01$; $a^{e}_{t-3}$ represent the attacker’s enemy unit moving westward, $a^{e}_{t-2}$ is moving southward, $a^{e}_{t-1}$ is moving eastward.
  • Trigger 3 $\mathcal{T}_{3}:=(\Psi_{3},\zeta_{3})$, where $\Psi_{3}(\mathbb{N}^{t}=5):=\psi_{(t-4,x)}\wedge\psi_{(t-4,y)}\wedge\psi_{(t-3,x)}\wedge\psi_{(t-3,y)}\wedge\psi_{(t-2,x)}\wedge\psi_{(t-2,y)}\wedge\psi_{(t-1,x)}\wedge\psi_{(t-1,y)}\wedge\psi_{(t,x)}\wedge\psi_{(t,y)}$, and $\zeta_{3}:=(a^{e}_{t-4},a^{e}_{t-3},a^{e}_{t-2},a^{e}_{t-1},a^{e}_{t})$, where $\psi_{(t-4,x)}:=x^{e}-x^{b}\equiv 2$, $\psi_{(t-4,y)}:=y^{e}-y^{b}\equiv 2$, $\psi_{(t-3,x)}:=x^{e}-x^{b}\equiv 1$, $\psi_{(t-3,y)}:=y^{e}-y^{b}\equiv 2$, $\psi_{(t-2,x)}:=x^{e}-x^{b}\equiv 1$, $\psi_{(t-2,y)}:=y^{e}-y^{b}\equiv 1$, $\psi_{(t-1,x)}:=x^{e}-x^{b}\equiv 2$, $\psi_{(t-1,y)}:=y^{e}-y^{b}\equiv 1$, $\psi_{(t,x)}:=x^{e}-x^{b}\equiv 2$, $\psi_{(t,y)}:=y^{e}-y^{b}\equiv 2$; $a^{e}_{t-4}$ represent a evader moving westward, $a^{e}_{t-3}$ is moving southward, $a^{e}_{t-2}$ is moving eastward, and $a^{e}_{t-1}$ is moving northward.
@@
In each poisoned episode, we search if there is an observable enemy unit (or evader) whose position relative to the unit controlled by the BLAST agent satisfies the first spatial constraint in $\Psi$. If found, we will take control of the enemy unit (or evader) to behave following $\zeta$ defined in the corresponding $\mathcal{T}$, and return it to its heuristic controller after the trigger. During BLAST model training, we set $\lambda=0.5$ in hacked reward, the size of both clean replay buffer $\mathcal{B}_{c}$ and poisoned replay buffer $\mathcal{B}_{p}$ to 5000, the size of batch to 32, the discount factor $\gamma=0.99$, and the learning rate $\alpha=5e-4$. For VDN and QMIX, we set the update frequency of the target network to 200 steps, and the greedy factor $\sigma=0.05$. For MAPPO, we set the GAE parameter $\epsilon=0.2$. For the 8m and 3m maps, we set the poisoning rate $p=0.05$, and the attack period $L=20$. For 2s3z map, we set $p=0.02$ and $L=40$. For the Pursuit environment, we set $p=0.05$ and $L=40$.

### v2

@@ Abstract
Recent studies have shown that cooperative multi-agent deep reinforcement learning (c-MADRL) is under the threat of backdoor attacks. Once a backdoor trigger is observed, it will perform malicious actions, resulting in failures or achieving malicious goals. However, existing backdoor attacks suffer from several issues, e.g., instant trigger patterns lack stealthiness, the backdoor is trained or activated by an additional network, or all agents are backdoored. To this end, in this paper, we propose a novel Backdoor Leverage Attack againST c-MADRL, BLAST, which attacks the entire multi-agent team by embedding the backdoor only in a single agent. Firstly, we introduce adversary spatiotemporal behavior patterns as the backdoor trigger rather than manually injected fixed visual patterns or instant status and control the period to perform malicious actions. This method can guarantee the stealthiness and practicality of BLAST. Secondly, we hack the original reward function of the backdoor agent via unilateral guidance to inject BLAST, to achieve the leverage attack effect that can pry open the entire multi-agent system via a single backdoor agent. We evaluate our BLAST against 3 classic c-MADRL algorithms (VDN, QMIX, and MAPPO) in 2 popular c-MADRL environments (SMAC and Pursuit), and 3 existing defense mechanisms. The experimental results demonstrate that BLAST can achieve a high attack success rate while maintaining a low clean performance variance rate.
@@ I Introduction
Cooperative multi-agent deep reinforcement learning (c-MADRL) is a significant branch of deep reinforcement learning (DRL). c-MADRL enables multiple agents to cooperate to achieve the same goal to solve complex tasks and has found applications in many areas, such as computation offloading [1, 2], cooperative game [3], and autonomous driving [4, 5]. However, current research [6, 7, 8, 9, 10, 11] has revealed that DRL is vulnerable to a serious threat known as backdoor attacks. This attack involves injecting specific triggers into the training dataset or process. Once a DRL policy is injected with a backdoor, it will behave in an unexpected or even malicious way when the trigger occurs. Since c-MADRL inherits the characteristics of DRL, it also suffers from this threat.
@@
Figure 1: A comparison of backdoor attacks against c-MADRL. (a) Most existing attacks use instant visual pattern triggers and require the implantation of backdoors into all agents to achieve the attack effect. (b) Our BLAST uses spatiotemporal behavior pattern triggers and has leverage effects that only require implanting the backdoor into a single agent, but can achieve the system-wide attack effect.
@@
As of now, only a handful of studies [12, 13, 14, 15] have delved into backdoor attacks against c-MADRL. Attacks [13, 12] directly implant backdoors in all agents and trigger them multiple times to carry out the attack, as shown in Fig. 1(a). While it is rather effortless to implant backdoor attacks in all agents, they come with a high injection cost and lack sufficient concealment. Attacks [14, 15] poison a single agent to attack the entire system, but ignore the mutual influence between agents [16]. This influence is the key to achieving fast cooperation in multi-agent systems [16] and can therefore weaken or even block the effectiveness of attacks [17]. In addition, existing attack methods all use specific instant visual patterns (shown in Fig. 1(a)) as triggers, e.g., visual modified map patterns [13], spectrum signals [12], and agent spatial distance [15]. Once the trigger is observed, they immediately execute instantaneous or continuous attack actions. However, this paired backdoor triggering and attack behavior can be easily detected by data anomaly detectors, e.g., [18, 19]. Hence, although these backdoor attacks are effective against c-MADRL, they are still far from perfect, which should possess two key characteristics: 1) good concealment and operability, and 2) minimal injection cost with a high impact.
@@
To achieve these two characteristics, we propose a novel stealthy backdoor leverage attack against c-MADRL, BLAST, which only requires implanting a backdoor in a single agent to pry an entire multi-agent team into failure, as shown in Fig. 1(b). We call this effect the leverage attack effect. First, most existing c-MADRL algorithms (e.g., VDN [20], QMIX [21], COMA [22], and MAPPO [23]) use recurrent neural networks (RNN) to memorize past information and make effective decisions in combination with current observations, thereby overcoming partial observability [24]. This, however, also allows attackers to hide backdoor triggers in unobservable states [5, 25]. Hence, to achieve the first characteristic, we distribute our BLAST’s trigger across a sequence of observations over a short period rather than within a single epoch. Once BLAST is deployed, the attacker only needs to act as a moving object (e.g., a user-controlled enemy unit in StarCraft) and perform a specific sequence of actions around the BLAST agent [^1] In the follow-up, we use this term to denote a c-MADRL agent that is implanted with our BLAST backdoor attack. in a short period following the trigger to activate the attack. Besides, we perform BLAST’s attack behaviors within a controllable period after the trigger occurs, rather than immediately. This method of decoupling attack triggers and actions and distributing them in different time series can ensure that BLAST achieves high concealment and a high attack success rate at a low poisoning rate.
@@
For the second characteristic, BLAST does not require additional networks to achieve the leverage attack effect. Specifically, c-MADRL agents are commonly trained by enlarging mutual influence between agents to achieve faster cooperation [16]. Under this influence, actions taken by agents will affect each other. Based on this, we design a reward hacking method for backdoor injection, in which we introduce the unilateral influence filter [17] to only enlarge the influence of the BLAST agent on other agents, but eliminate the opposite impact. It consists of two reward items: the former encourages the BLAST agent to perform actions that have long-term malicious effects on clean agents through target failure state guidance; the latter incentivizes the BLAST agent to quickly find actions that can mislead its teammates into performing non-optimal actions. With this hacking method, BLAST can quickly attack the entire system by poisoning only one agent.
@@
  • We use spatiotemporal behavior patterns as triggers of our BLAST, and its attack period is controllable, which enhances its concealment and operability;
@@ II Background
Most popular c-MADRL algorithms adopt the centralized training and decentralized execution (CTDE) framework [27, 28]. The CTDE framework centralizes the evaluation of the joint policies of all agents during training to address the issue of environmental non-stationarity. During execution, each agent is dispersed and executed separately, relying solely on its own local observation and observation history for action selection, greatly alleviating the problem of low efficiency in policy execution. c-MADRL can be roughly divided into two categories: value function decomposition (VFD) and centralized value function (CVF).
@@
CVF-based c-MADRL typically addresses the scalability and environmental non-stationarity issues by learning a centralized critic network and an independent actor network for each agent. MAPPO [23] is a CVF-based algorithm that applies the PPO algorithm [29] to multi-agent environments. Through methods such as value function normalization, action masking, and importance sampling, it effectively solves the problem of low data sample sampling efficiency in on-policy algorithms under limited computing resources. The objective function of MAPPO is as follows:
@@ III-B Attacker’s Capacities and Goals
Attacker’s Goals. With the above capabilities and limitations, the attacker’s main goal is to attack the entire multi-agent team by triggering the backdoor in a single agent during execution. Moreover, the implanted backdoor needs to be effective and stealthy. Specifically, if the backdoor trigger is present, the BLAST agent is capable of behaving maliciously or abnormally to disrupt the entire team; otherwise, it is capable of behaving normally like a clean agent. Besides, the backdoor triggers should be concealed, occur as infrequently as possible, and have a low poisoning rate.
@@ IV-A Spatiotemporal Backdoor Trigger
IV-B Reward Hacking based on Unilateral Influence
In c-MADRL systems, all agents typically interact with each other [31]; thus, there is a mutual influence relationship between agents. Due to this mutual influence, the sphere of influence of reward hacking methods in the existing backdoor attacks [7, 8, 32, 10] may be limited only to the backdoor agent itself rather than the entire system during the attack period. Therefore, we design a new reward hacking method based on the unilateral influence filter [17], which can eliminate the detrimental influence from other agents on the backdoor agent and only enable the influence from the backdoor agent to other agents.
@@ IV-B Reward Hacking based on Unilateral Influence
To enhance the effectiveness of BLAST, we further consider the perspective of action deviation, where the BLAST agent performs a malicious action, causing the clean agents’ actions at the next time step to deviate from the original optimal actions. We set the hacked reward term based on action deviation as follows:
@@ IV-C BLAST Model Training
Firstly, the attacker specifies a spatiotemporal behavior trigger $\mathcal{T}:=(\Psi,\zeta)$ with a trigger period $\mathbb{N}^{t}$. Besides, an attack period $L$ is introduced to attack only $L$ time steps after the trigger appears completely, which can enhance the stealthiness of the backdoor attacks. In each training episode, the attacker determines whether to poison with the poisoning rate $p$. If to poison, the attacker will insert the backdoor trigger (Line 3). At each time step $t$ in an episode, clean agents select their actions according to their policies, and the BLAST agent chooses a random action with the probability $\sigma$, otherwise, it chooses an action according to its policy $\pi^{b}$ (Line 5-7). During the attack period $L$, the attacker will hack the reward $r_{t}$ as Equ (5) (Line 11-14). To ensure the effective training of the BLAST model, we set up two replay buffers, $\mathcal{B}_{p}$ and $\mathcal{B}_{c}$, to store poisoned and clean episodes, respectively (Line 17-21). To update the BLAST model, we randomly sample a batch of episode data from replay buffer $\mathcal{B}_{p}$ or $\mathcal{B}_{c}$ with probability $p$ or $1-p$ (Line 22-23).
@@ V-A1 Environments
We use StarCraft Multi-Agent Challenge (SMAC) [3] and Pursuit [33] in pettingzoo [34] as experimental environments, as shown in Fig. 3: (a) SMAC is an environment of two teams against each other where all allied units (i.e., the red side) are controlled by c-MADRL agents, while enemy units (i.e., the blue side) are controlled by users (including a attacker). If the enemies’ health decreases, the agent team will receive corresponding positive rewards; if an enemy unit dies, the agent team will receive a reward of 10; if all enemy units die (i.e. win), the agent team will receive a reward of 200; if all allied units die (i.e. fail), the agent team will receive a reward of 0. We choose 8m, 3m, and 2s3z as our test maps, with 8, 3, and 5 agents in the agent team, respectively. (b) In the Pursuit environment, the pursuers are controlled by agents. Every time the agents fully surround an evader, each of the surrounding agents receives a reward of 5, and the evader is removed from the environment. Agents also receive a reward of 0.01 every time they touch an evader. We set the number of agents to 8 and the number of evaders to 10. Each agent receives a fixed penalty of -0.01 per time step, which means that the longer it takes to catch the same number of evaders, the lower the team reward.
@@ V-A3 Backdoor Trigger and Attack Settings
We use the Trigger-1 $\mathcal{T}_{1}:=(\Psi_{1},\zeta_{1})$ defined in Example 1 to evaluate our backdoor attacks in the 8m and 2s3z map. In the 3m map and Pursuit environment, we use Trigger-2 and Trigger-3 defined as follows, respectively:
  • Trigger-2 $\mathcal{T}_{2}:=(\Psi_{2},\zeta_{2})$, where $\Psi_{2}(\mathbb{N}^{t}=4):=\wedge\psi_{(t-3,x)}\wedge\psi_{(t-3,y)}\wedge\psi_{(t-2,x)}\wedge\psi_{(t-2,y)}\wedge\psi_{(t-1,x)}\wedge\psi_{(t-1,y)}\wedge\psi_{(t,x)}\wedge\psi_{(t,y)}$, and $\zeta_{2}:=(a^{e}_{t-3},a^{e}_{t-2},a^{e}_{t-1},a^{e}_{t})$, where $\psi_{(t-3,x)}:=0.98<x^{e}-x^{b}<1.00$, $\psi_{(t-3,y)}:=-0.10<y^{e}-y^{b}<0.10$, $\psi_{(t-2,x)}:=0.60<x^{e}-x^{b}<0.62$, $\psi_{(t-2,y)}:=0.60<y^{e}-y^{b}<0.62$, $\psi_{(t-1,x)}:=0.60<x^{e}-x^{b}<0.62$, $\psi_{(t-1,y)}:=-0.62<y^{e}-y^{b}<-0.60$, $\psi_{(t,x)}:=0.98<x^{e}-x^{b}<1.00$, $\psi_{(t,y)}:=-0.01<y^{e}-y^{b}<0.01$; $a^{e}_{t-3}$ represent the attacker’s enemy unit moving westward, $a^{e}_{t-2}$ is moving southward, $a^{e}_{t-1}$ is moving eastward.
  • Trigger-3 $\mathcal{T}_{3}:=(\Psi_{3},\zeta_{3})$, where $\Psi_{3}(\mathbb{N}^{t}=5):=\psi_{(t-4,x)}\wedge\psi_{(t-4,y)}\wedge\psi_{(t-3,x)}\wedge\psi_{(t-3,y)}\wedge\psi_{(t-2,x)}\wedge\psi_{(t-2,y)}\wedge\psi_{(t-1,x)}\wedge\psi_{(t-1,y)}\wedge\psi_{(t,x)}\wedge\psi_{(t,y)}$, and $\zeta_{3}:=(a^{e}_{t-4},a^{e}_{t-3},a^{e}_{t-2},a^{e}_{t-1},a^{e}_{t})$, where $\psi_{(t-4,x)}:=x^{e}-x^{b}\equiv 2$, $\psi_{(t-4,y)}:=y^{e}-y^{b}\equiv 2$, $\psi_{(t-3,x)}:=x^{e}-x^{b}\equiv 1$, $\psi_{(t-3,y)}:=y^{e}-y^{b}\equiv 2$, $\psi_{(t-2,x)}:=x^{e}-x^{b}\equiv 1$, $\psi_{(t-2,y)}:=y^{e}-y^{b}\equiv 1$, $\psi_{(t-1,x)}:=x^{e}-x^{b}\equiv 2$, $\psi_{(t-1,y)}:=y^{e}-y^{b}\equiv 1$, $\psi_{(t,x)}:=x^{e}-x^{b}\equiv 2$, $\psi_{(t,y)}:=y^{e}-y^{b}\equiv 2$; $a^{e}_{t-4}$ represent a evader moving westward, $a^{e}_{t-3}$ is moving southward, $a^{e}_{t-2}$ is moving eastward, and $a^{e}_{t-1}$ is moving northward.
@@
In each poisoned episode, we search if there is an observable enemy unit (or evader) whose position relative to the unit controlled by the BLAST agent satisfies the first spatial constraint in $\Psi$. If found, we will take control of the enemy unit (or evader) to behave following $\zeta$ defined in the corresponding $\mathcal{T}$, and return it to its heuristic controller after the trigger. During BLAST model training, we set $\lambda=0.5$ in hacked reward, the size of both clean replay buffer $\mathcal{B}_{c}$ and poisoned replay buffer $\mathcal{B}_{p}$ to 5000, the size of batch to 32, the discount factor $\gamma=0.99$, and the learning rate $\alpha=5e^{-4}$. For VDN and QMIX, we set the update frequency of the target network to 200 steps, and the greedy factor $\sigma=0.05$. For MAPPO, we set the GAE parameter $\epsilon=0.2$. For the 8m and 3m maps, we set the poisoning rate $p=0.05$, and the attack period $L=20$. For 2s3z map, we set $p=0.02$ and $L=40$. For the Pursuit environment, we set $p=0.05$ and $L=40$.